myininaya (myininaya):

anyone besides satellite i want to see how to factor quadratics over real numbers ( not just the integers) without using the quadratic formula or without using completing the square

myininaya (myininaya):

an easy quadratic

myininaya (myininaya):

http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e3a5e000b8bf47d066015a8 here are some examples here

OpenStudy (anonymous):

Impressive...pure mathematician in the making:-)

myininaya (myininaya):

\[ax^2+bx+c\] you want to find two factors of a*c that have product a*c and have sum b so we have b=(b/2+z)+(b/2-z) a*c=(b/2+z)(b/2-z)=b^2/4-z^2 <- solve this for z

myininaya (myininaya):

then replace bx with (b/2+z)x+(b/2-z)x then factor by grouping :)

myininaya (myininaya):

thanks estudier :)

OpenStudy (anonymous):

U shouldn't give medal, no-one will come to the post:-) Delete it and post it again.

myininaya (myininaya):

delete this thread?

OpenStudy (anonymous):

You can copy the replies into the new one, don't give any medals, then people will come to look.

OpenStudy (anonymous):

People should see your method...

myininaya (myininaya):

lol estudier copy the replies from the link above into this link? i'm sorry im a retard

OpenStudy (anonymous):

I will do it for u (will be fun, I hope..)

myininaya (myininaya):

lol thanks

myininaya (myininaya):

thanks estudier your awesome sorry im slow